If a tank with 4.2kg of water is leaking at a rate of 0.0135kg/s. How many hours will it take for the tank to empty
1
Expert's answer
2020-09-14T08:25:39-0400

time it will take to empty vessel = 4.2 / 0.0135 = 311.11 seconds

Now time taken in hour = 311.11/3600 =0.0864 = 0.09 hour